Divisional Performance Analysis and EvaluationDivisional Performance Analysis and Evaluation The vice president of operations of Recycling Industries is evaluating the per2. Using the DuPont formula for return on investment, determine the profit margin, investment turnover, and return on investm
Answer #1

divisional income statement

particular

business division

consumer division

sale

4,28,00,000

5,60,00,000

cost of goods sold

2,35,00,000

3,05,00,000

gross profit

1,93,00,000

2,55,00,000

operating expense

1,14,24,800

1,43,00,000

operating income

78,75,200

1,12,00,000

business division

profit margin = net profit / sale x 100

                            = 78,75,200 / 4,28,00,000 x 100

                            = 18.4%

investment turnover = income / invested assets

                                            = 78,75,200 / 34,240,000

                                            = 0.23

ROI = Return / investment x 100

        = 78,75,200 / 34,240,000 x 100

        = 23%

consumer division

profit margin = net profit / sale x 100

                            = 1,12,00,000 / 5,60,00,000 x 100

                            = 20%

investment turnover = income / invested assets

                                            = 1,12,00,000 / 70,000,000

                                            = 0.16

ROI = Return / investment x 100

        = 1,12,00,000 / 70,000,000 x 100

        = 16%

division

profit margin

investment turnover

ROI

business

18.40%

0.23

23%

consumer

20%

0.16

16%

business division

residual income = operating income -(minimum required return x invested assets)

                                  = 78,75,200 - (10% x 34,240,000)

                                  = 78,75,200 - 3,424,000

                                  = 4,451,200

consumer division

residual income = operating income -(minimum required return x invested assets)

                                  = 1,12,00,000 - (10% x 70,000,000)

                                  = 1,12,00,000 - 7,000,000

                                  = 4,200,000

on the basis of net operating income consumer division is more profitable. on the basis of residual income business division is more profitable